# Riemannian Geometry of Optimal Rebalancing in Dynamic Weight Automated Market Makers

### 09 Mar 2026 v2

@@ 4 SLERP on the sphere under Hellinger embedding
Among all $f$-step paths $\boldsymbol{\mathbf{w}}(t_{0}),\boldsymbol{\mathbf{w}}(t_{1}),\ldots,\boldsymbol{\mathbf{w}}(t_{f})$ on the simplex from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$, the one that minimises the total quadratic loss $\sum_{k=1}^{f}\sum_{i}(\Delta w_{i}^{(k)})^{2}/(2\,w_{i}^{(k-1)})$ is the constant-speed geodesic of the Fisher–Rao metric: SLERP in Hellinger coordinates. Each step incurs equal loss $2\Omega^{2}/f^{2}$, where $\Omega=\arccos\!\left(\sum_{i=1}^{N}\sqrt{w_{i}^{\mathrm{start}}\,w_{i}^{\mathrm{end}}}\right)$, giving total loss $2\Omega^{2}/f$.
@@
The quadratic loss per step is the squared arc-length under $\tfrac{1}{2}g$. By convexity of $x\mapsto x^{2}$, the total loss $\sum_{k=1}^{f}(\mathrm{d}s_{k})^{2}\geq S^{2}/f$ for any path of total arc-length $S$, with equality if and only if all steps are equal (Appendix B). The minimum over paths is attained by the geodesic (shortest $S$). Under the Hellinger embedding (§3), the metric is the scaled round metric on $S^{N-1}_{+}$, whose geodesics are great circles parameterised at constant angular velocity, that is: SLERP [3]. ∎
4.1 The SLERP formula
@@
$\mathcal{L}_{k}=(\mathrm{d}s_{k})^{2}=\frac{2\Omega^{2}}{f^{2}}$ (13)
@@ 4.2 Two-token case
$w(k)=\sin^{2}\!\left(\theta^{\mathrm{start}}+\frac{k}{f}\left(\theta^{\mathrm{end}}-\theta^{\mathrm{start}}\right)\right),$ (14)
@@
5 SLERP midpoint equals (AM+GM)/normalise midpoint
Does the prior (AM+GM)/normalise heuristic [1], Eq (5), connect to the SLERP geodesics? The midpoints of the two trajectories are identical.
@@ 5 SLERP midpoint equals (AM+GM)/normalise midpoint
For any NN-token pool with start weights $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ and end weights $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$, the SLERP midpoint ($\boldsymbol{\eta}(t)\leftarrow\dfrac{\sin\!\left((1-t)\,\Omega\right)}{\sin\Omega}\,\boldsymbol{\eta}^{\mathrm{start}}+\dfrac{\sin\!\left(t\,\Omega\right)}{\sin\Omega}\,\boldsymbol{\eta}^{\mathrm{end}}$ at $t=k/f=1/2$) is exactly equal to the (AM+GM)/normalise midpoint of [1].
@@
$w_{i}^{\mathrm{SLERP}}=\frac{\left(\sqrt{w_{i}^{\mathrm{start}}}+\sqrt{w_{i}^{\mathrm{end}}}\right)^{2}}{\sum_{j}\left(\sqrt{w_{j}^{\mathrm{start}}}+\sqrt{w_{j}^{\mathrm{end}}}\right)^{2}},$ (15)
@@
$\left(\sqrt{w_{i}^{\mathrm{start}}}+\sqrt{w_{i}^{\mathrm{end}}}\right)^{2}=2\!\left(\underbrace{\frac{w_{i}^{\mathrm{start}}+w_{i}^{\mathrm{end}}}{2}}_{\mathrm{AM}_{i}}+\underbrace{\sqrt{w_{i}^{\mathrm{start}}\,w_{i}^{\mathrm{end}}}}_{\mathrm{GM}_{i}}\right).$ (16)
@@
$w_{i}^{\mathrm{SLERP}}=\frac{\mathrm{AM}_{i}+\mathrm{GM}_{i}}{\sum_{j}\left(\mathrm{AM}_{j}+\mathrm{GM}_{j}\right)}=\breve{w}_{i},$ (17)
which is precisely the (AM+GM)/normalise formula (Eq (5), from [1] at $t=1/2$). ∎
@@
Figure 3: Visual proof that the SLERP midpoint equals (AM+GM)/normalise (in the N=3 case). To return to weight space we undo the Hellinger embedding $\eta_{i}=\sqrt{w_{i}}$, i.e. square. Each square decomposes $(\sqrt{w_{i}^{\mathrm{s}}}+\sqrt{w_{i}^{\mathrm{e}}})^{2}=(w_{i}^{\mathrm{s}}+w_{i}^{\mathrm{e}})+2\sqrt{w_{i}^{\mathrm{s}}\,w_{i}^{\mathrm{e}}}=2(\mathrm{AM}_{i}+\mathrm{GM}_{i})$. Red diagonal blocks contribute $2\,\mathrm{AM}_{i}$; blue off-diagonal blocks contribute $2\,\mathrm{GM}_{i}$. The factor of 2 cancels with normalisation, giving us the (AM+GM)/normalise heuristic exactly.
@@
This means an on-chain implementation needs only addition, multiplication, square root, and normalisation, so no $\arccos$ or $\sin$ evaluations, to produce the optimal interpolation at power-of-two step counts.
@@
Figure 4: Recursive bisection on the positive orthant of $S^{N-1}_{+}$. At each depth $d$, the (AM+GM)/normalise midpoint is inserted between every pair of adjacent points (bright green: newly inserted; dark green: inherited from previous depth). By depth 3, the $2^{d}+1=9$ points densely trace the geodesic arc, all computed without trigonometric functions.
5.1 Comparison at general $t$ and with Lambert W
@@
$w_{i}^{\mathrm{SLERP}}(t)=\left[\frac{\sin\!\left((1-t)\,\Omega\right)}{\sin\Omega}\sqrt{w_{i}^{\mathrm{start}}}+\frac{\sin\!\left(t\,\Omega\right)}{\sin\Omega}\sqrt{w_{i}^{\mathrm{end}}}\right]^{\!2},$ (18)
@@
Table 1: Agreement between midpoint formulae in powers of $u_{i}=\Delta w_{i}/w_{i}$.
Pair of methods Midpoint General $t$
SLERP vs (AM+GM)/normalise Exact Differ at $O(u^{3})$
SLERP vs Lambert W (normalised) Differ at $O(u^{3})$ Differ at $O(u^{3})$
(AM+GM)/normalise vs Lambert W (normalised) Differ at $O(u^{3})$ Differ at $O(u^{3})$
6 Bounding SLERP’s sub-optimality on the exact KL cost
SLERP minimises the quadratic approximation to the per-step KL cost (Corollary 2). The following theorem bounds the gap to the global optimum on the exact cost.
@@ 6 Bounding SLERP’s sub-optimality on the exact KL cost
$0\;\leq\;C_{\mathrm{SLERP}}(f)-C_{*}(f)\;\leq\;\frac{A\,\Omega^{4}}{f^{3}},$ (19)
where AA depends on the weight configuration and $\epsilon$ but not on $f$. The relative sub-optimality is $O(\Omega^{2}/f^{2})$.
@@
Deferred to Appendix C. The explicit constant is $A=128\,N/\epsilon^{4}$. ∎
@@
Expanding the per-step KL cost beyond the quadratic approximation that SLERP minimises (Appendix C), the leading remainder is $R_{k}=-\frac{1}{6}\sum_{i}w_{i}(u_{i}^{(k)})^{3}+O(u^{4})$, which encodes the asymmetry of the KL divergence: it is negative when weights increase ($u_{i}>0$, so the true KL cost is less than the quadratic) and positive when weights decrease. SLERP, as the geodesic of the symmetric Fisher–Rao metric ($\alpha=0$), treats both directions equally and does not exploit this asymmetry. The exact KL-optimal path would take slightly larger steps where weights increase and slightly smaller steps where they decrease, but the available improvement is only $O(u^{2})$ per step, hence the bound.

### 31 Mar 2026 v3

@@ 3.2. SLERP on the sphere under Hellinger embedding
Among all $f$-step paths on the simplex from $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ to $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$, the path minimising the sum of squared Fisher–Rao arc-lengths $\sum_{k=1}^{f}(\mathrm{d}s_{k})^{2}$ is the constant-speed geodesic: SLERP in Hellinger coordinates, where the geodesic angle is $\Omega=\arccos(\mathrm{BC})$ and $\mathrm{BC}=\sum_{i}\sqrt{w_{i}^{s}w_{i}^{e}}$ is the Bhattacharyya coefficient [Bhattacharyya, 1943].
@@
Under the Hellinger embedding (§3.1), the Fisher–Rao metric is $4\times$ the round metric on $S^{N-1}_{+}$, so $\mathrm{d}s_{k}=2\alpha_{k}$ where $\alpha_{k}$ is the geodesic angle subtended by step $k$. By convexity of $x\mapsto x^{2}$, $\sum_{k=1}^{f}\alpha_{k}^{2}\geq(\sum_{k}\alpha_{k})^{2}/f$ with equality iff all $\alpha_{k}$ are equal (Appendix B). The minimum over paths is attained by the geodesic (shortest total angle $\sum_{k}\alpha_{k}=\Omega$), traversed at constant angular speed ($\alpha_{k}=\Omega/f$): that is, SLERP [Shoemake, 1985]. ∎
Along the SLERP path, each step subtends equal arc $\Omega/f$, giving per-step loss $\mathrm{d}s_{k}^{2}/2=2\Omega^{2}/f^{2}$ and total loss $2\Omega^{2}/f$. The per-step loss equals the quadratic loss kernel of Corollary 3 at leading order. [^4] The base-point quadratic form $\sum_{i}(\Delta w_{i})^{2}/(2w_{i})$ differs from the exact $\mathrm{d}s^{2}/2$ by $O(\|\Delta\boldsymbol{\mathbf{\eta}}\|^{3})$ per step, giving a total discrepancy of $O(\Omega^{3}/f^{2})$.
3.3. The SLERP formula
@@
(13) $\mathcal{L}_{k}=(\mathrm{d}s_{k})^{2}=\frac{2\Omega^{2}}{f^{2}}$
@@ 3.4. Two-token case
(14) $w(k)=\sin^{2}\!\left(\theta^{\mathrm{start}}+\frac{k}{f}\left(\theta^{\mathrm{end}}-\theta^{\mathrm{start}}\right)\right),$
@@
3.5. SLERP midpoint equals (AM+GM)/normalise midpoint
Does the (AM+GM)/normalise heuristic [Willetts and Harrington, 2024a], Eq (5), connect to the SLERP geodesics? The midpoints of the two trajectories are identical.
@@ 3.5. SLERP midpoint equals (AM+GM)/normalise midpoint
For any $N$-token pool with start weights $\boldsymbol{\mathbf{w}}^{\mathrm{start}}$ and end weights $\boldsymbol{\mathbf{w}}^{\mathrm{end}}$, the SLERP midpoint
$\boldsymbol{\eta}(t)=\frac{\sin\!\bigl((1-t)\,\Omega\bigr)}{\sin\Omega}\,\boldsymbol{\eta}^{\mathrm{start}}+\frac{\sin\!\bigl(t\,\Omega\bigr)}{\sin\Omega}\,\boldsymbol{\eta}^{\mathrm{end}}$
evaluated at $t=k/f=1/2$ is exactly equal to the (AM+GM)/normalise midpoint of [Willetts and Harrington, 2024a].
@@
(15) $w_{i}^{\mathrm{SLERP}}=\frac{\left(\sqrt{w_{i}^{\mathrm{start}}}+\sqrt{w_{i}^{\mathrm{end}}}\right)^{2}}{\sum_{j}\left(\sqrt{w_{j}^{\mathrm{start}}}+\sqrt{w_{j}^{\mathrm{end}}}\right)^{2}},$
@@
(16) $\left(\sqrt{w_{i}^{\mathrm{start}}}+\sqrt{w_{i}^{\mathrm{end}}}\right)^{2}=2\!\left(\underbrace{\frac{w_{i}^{\mathrm{start}}+w_{i}^{\mathrm{end}}}{2}}_{\mathrm{AM}_{i}}+\underbrace{\sqrt{w_{i}^{\mathrm{start}}\,w_{i}^{\mathrm{end}}}}_{\mathrm{GM}_{i}}\right).$
@@
(17) $w_{i}^{\mathrm{SLERP}}=\frac{\mathrm{AM}_{i}+\mathrm{GM}_{i}}{\sum_{j}\left(\mathrm{AM}_{j}+\mathrm{GM}_{j}\right)}=\breve{w}_{i},$
which is precisely the (AM+GM)/normalise formula (Eq (5), from [Willetts and Harrington, 2024a] at $t=1/2$). ∎
@@
Figure 3. Visual proof that the SLERP midpoint equals (AM+GM)/normalise (in the N=3 case). To return to weight space we undo the Hellinger embedding $\eta_{i}=\sqrt{w_{i}}$, i.e. square. Each square decomposes: $(\!\sqrt{w_{i}^{\mathrm{s}}}+\sqrt{w_{i}^{\mathrm{e}}})^{2}=(w_{i}^{\mathrm{s}}+w_{i}^{\mathrm{e}})+2\!\sqrt{w_{i}^{\mathrm{s}}\,w_{i}^{\mathrm{e}}}$ $2(\mathrm{AM}_{i}+\mathrm{GM}_{i})$. Red diagonal blocks contribute $2\,\mathrm{AM}_{i}$; blue off-diagonal blocks contribute $2\,\mathrm{GM}_{i}$. The factor of 2 cancels with normalisation, giving us the (AM+GM)/normalise heuristic exactly.
3.6. Trig-free recursive bisection
@@
This means an on-chain implementation needs only addition, multiplication, square root, and normalisation (no $\arccos$ or $\sin$ evaluations) to produce the optimal interpolation at power-of-two step counts.
@@
Figure 4. Recursive bisection on the positive orthant of $S^{N-1}_{+}$. At each depth $d$, the (AM+GM)/normalise midpoint is inserted between every pair of adjacent points (bright green: newly inserted; dark green: inherited from previous depth). By depth 3, the $2^{d}+1=9$ points densely trace the geodesic arc, all computed without trigonometric functions.
3.7. Comparison at general $t$ and with Lambert W
@@
(18) $w_{i}^{\mathrm{SLERP}}(t)=\left[\frac{\sin\!\left((1-t)\,\Omega\right)}{\sin\Omega}\sqrt{w_{i}^{\mathrm{start}}}+\frac{\sin\!\left(t\,\Omega\right)}{\sin\Omega}\sqrt{w_{i}^{\mathrm{end}}}\right]^{\!2},$
@@
Table 1. Agreement between midpoint formulae in powers of $u_{i}=\Delta w_{i}/w_{i}^{\mathrm{start}}$.
Pair of methods Midpoint General $t$ Source of $O(u^{3})$ difference
SLERP vs (AM+GM)/normalise Exact $O(u^{3})$ SLERP coefficients vs $(1{+}u)^{t}$
SLERP vs Lambert W (normalised) $O(u^{3})$ $O(u^{3})$ Quadratic vs exact KL objective
(AM+GM)/normalise vs Lambert W (normalised) $O(u^{3})$ $O(u^{3})$ Quadratic vs exact KL objective
3.8. Sub-optimality of SLERP on the exact KL cost
SLERP minimises the quadratic approximation to the per-step KL cost (Corollary 5). Theorem 9 bounds the gap to the global optimum on the exact cost.
@@ 3.8. Sub-optimality of SLERP on the exact KL cost
(19) $0\;\leq\;C_{\mathrm{SLERP}}(f)-C_{*}(f)\;\leq\;\frac{A\,\Omega^{4}}{f^{3}},$
where $A$ depends on the weight configuration and $\epsilon$ but not on $f$. The relative sub-optimality is $O(\Omega^{2}/f^{2})$.
@@
Deferred to Appendix C. The explicit constant is $A=86\,N/\epsilon^{4}$. ∎
The bound is qualitatively informative (rate in $f$ and $\Omega$) rather than quantitatively tight: the constant $A$ is conservative, and the numerics show much smaller gaps than the bound predicts. For the $N=3$ setup ($\epsilon=0.05$), the condition $f\geq 4\Omega/\epsilon\approx 42$ is easily satisfied at any practical step count.
@@
Expanding the per-step KL cost beyond the quadratic approximation that SLERP minimises (Appendix C), the leading remainder is $R_{k}=-\frac{1}{6}\sum_{i}w_{i}(u_{i}^{(k)})^{3}+O(u^{4})$, which encodes the asymmetry of the KL divergence: it is negative when weights increase ($u_{i}>0$, so the true KL cost is less than the quadratic) and positive when weights decrease. SLERP is the geodesic of the symmetric Fisher–Rao metric ($\alpha=0$) and ignores this asymmetry entirely. The exact KL-optimal path would take slightly larger steps where weights increase and slightly smaller ones where they decrease. But the per-step gain is only $O(u^{2})$, accumulating to $O(\Omega^{4}/f^{3})$ (Theorem 9). A $[1,1]$ Padé approximant resums this cubic remainder into a rational function matching the exact KL cost through $O(u^{4})$ (Appendix M).